You're riding a unicorn at 25m/s and come to a uniform stop at a red light 20m away. What's your acceleration ?
1
Expert's answer
2018-02-26T09:08:07-0500
We can find acceleration of the unicorn from the kinematic equation:
v^2=v_0^2+2as,
here, v_0=25 m⁄s is the initial velocity of the unicorn, v=0 is the final velocity of the unicorn when it stops at the red light, a is the acceleration of the unicorn and s=20 m is the distance from the red light at which the unicorn begins to stop.
Then, we get:
a=-(v_0^2)/2s=-(25 m/s)^2/(2∙20 m)=-15.6 m/s^2 .
The sign minus indicates that the unicorn is decelerates.
Answer:
a=-15.6 m/s^2 .